Ms. Gonzalez would like to paint 4 of her classroom tables black. She knows that she can cover 50 square feet of an area with 1 quart of paint. She measured the top of the table to be 12 feet by 3 feet and the height to be 2 inches. If Ms. Gonzalez is only concerned with painting the top of the 4 table and the 4 sides of the table, how many quarts of paint does Ms. Gonzalez need to buy?
1
Expert's answer
2012-09-25T11:12:11-0400
The area of the table needed to be painted is

A = 12[ft]*3[ft] + 2*12[ft]*2[ft] + 2*3[ft]*2[ft] = 96 ft².

As there are four tables, the total area needed to be painted is

A = 4*96[ft²] = 384 ft².

So, Ms. Gonzalez need to buy

N = 384/50 = 7.68

quarts of paint.